### v2

# B-VLLM: A Vision Large Language Model with Balanced Spatio-Temporal Tokens

Abstract
Recently, Vision Large Language Models (VLLMs) with integrated vision encoders have shown promising performance in vision understanding. They encode visual content into sequences of visual tokens, enabling joint processing of visual and textual data. However, understanding videos, especially long videos, remains a challenge as the rapid growth of visual tokens during video encoding risks exceeding VLLMs’ context window length and significantly escalates computational cost. To restrict the number of visual tokens, existing VLLMs either: (1) uniformly downsample videos into a fixed number of frames or (2) reducing the number of visual tokens encoded from each frame. We argue that the former neglects temporal dynamics in videos, while the latter fails to preserve spatial details within individual frame. In this work, we propose Balanced-VLLM (B-VLLM), a novel VLLM framework designed to model task relevant spatio-temporal cues, while restricting the number of visual tokens within the VLLM’s context window length. Central to our framework is a text-conditioned adaptive frame selection module that dynamically identifies task-relevant frames, which are further de-duplicated with a temporal frame token merging strategy. The visual tokens of these frames then undergo spatial token sampling and an optional spatial token merging strategy for granular control against the token budget. Experiments demonstrate the effectiveness of B-VLLM in balancing the number of frames and visual tokens, moreover, our proposed method introduces 10% performance gain on MVBench. Code is available at https://github.com/zhuqiangLu/B-VLLM.git.
22footnotetext: Corresponding authors.
1 Introduction
(a) Context window limit.
(b) Key-hints missed in sampling.
(c) Video LLM performance on image-based spatial-only benchmarks.
Figure 1: Illustration of the limitations of existing VLLMs.
Large Language Models (LLMs) have demonstrated remarkable capability and generalizability in natural language understanding [4, 56, 51, 11]. Recently, there have been several notable efforts to extend LLMs into Vision Large Language Models (VLLMs) for visual understanding [2, 33, 62]. These efforts primarily aim to adapt LLMs’ contextual reasoning capabilities to multimodal settings like video captioning and visual question answering, enabling them to generate accurate textual responses based on visual and textual context provided by users. Unlike the extension of vision models to downstream tasks [19, 57, 37, 61, 8], bridging the modality gap between textual and visual contexts poses significant challenges [33]. Existing VLLMs typically use pretrained Vision Transformers (ViTs) [10] to encode images into sequences of visual tokens. These tokens are subsequently integrated with text token sequences as input to the VLLMs [33, 32]. This approach compresses spatial contexts from images into 1D sequences and has demonstrated its effectiveness across various computer vision tasks [14]. However, these VLLMs are primarily designed for single-image scenarios. As illustrated in Figure 1(a), extending them to video data often compromises performance, as the number of visual tokens exceeds the length limit of the VLLM’s context window.
To address this issue, methods like LLaMA-VID [30], MovieChat [50] and LongVLM [54] reduce the number of visual tokens to encode each frame. Alternatively, methods such as VideoLLaMA2 [9], VideoLLaMA [58] and VideoChat2 [27] sample a fixed number of frames uniformly from a video to prevent an overload of visual tokens. While being effective in controlling the number of visual tokens for video processing, these methods struggle to balance spatio-temporal cues. As depicted in Figure 1(b), VLLMs using uniform frame sampling tend to underperform on temporally relevant tasks. As shown in Figure 1(c), methods that reduce the number of visual tokens per frame often perform poorly on spatially relevant tasks.
We argue that this is due to a spatio-temporal token imbalance: reducing frame-level visual tokens causes temporal cues to dominate, whereas uniform frame sampling leads to spatial cues to be overshadowed. To address these issues, we propose Balanced-VLLM (B-VLLM), a novel VLLM framework to balance spatio-temporal cues in video data conditioned by task-related text prompts. B-VLLM processes each frame with both coarse-grained and fine-grained visual tokens. A learnable frame selection module leverages coarse-grained tokens to identify the frames most relevant to the video understanding task, guided by the text prompt. The fine-grained visual tokens of these selected frames are then processed by a spatial visual token sampling module to retrieve those most task-relevant ones. For granular control over the final number of generated visual tokens, we introduce an iterative spatial token merging strategy that averages similar visual tokens until the expected number of tokens is reached. Finally, the resulting visual tokens are mapped into the LLM feature space and integrated with the sequence of textual tokens encoded from the text prompt before being fed into the LLM. Our proposed method achieves SOTA performance on various benchmarks. Notably, our method introduces up to 10% performance improvement on MVBench [27]
The main contributions are summarized as follows:
  • We propose B-VLLM, a novel Vision LLM framework that dynamically balances spatio-temporal tokens for video understanding, especially for long videos.
  • We devise an adaptive token selection strategy by leveraging both coarse-grained and fine-grained tokens.
  • We undertake comprehensive experiments to show that B-VLLM is effective in different VLLM architectures for video understanding.
  • We demonstrate that the B-VLLM framework can be generalized to different datasets and VLLM settings.
2 Related Work
2.1 Vision Large Language Models
Vision Large Language Models (VLLMs) extend existing Large Language Models (LLMs) to the domain of visual understanding through visual-instruction tuning [33]. Visual content is encoded as sequences of visual tokens using a pretrained CLIP [44] vision encoder. These visual tokens are then projected into the LLM hyperspace through a trainable MLP layer [33, 30, 9, 31]. However, in multi-image or video scenarios, the number of vision tokens often exceeds the VLLM’s context window limit, leading to performance degrading and high computation costs. To address this, Open-Flamingo[2], MovieChat [50] and BLIP2  [26] reduce the number of vision tokens extracted from each frame by using transformer-based token selection networks, such as Q-Former [26] and Perceptive Resampler [2]. However, the number of vision tokens is proportional to the number of input images or frames, creating the risk of an overwhelming number of tokens when processing long videos. LLaMA-VID [30] addresses this by encoding each frame into only two visual tokens to enable hour-long video processing; however, this extreme compression rate limits LLaMA-VID’s ability to capture finer spatial details in each frame. Another common approach to controlling the number of visual tokens is to uniformly sample frames from a video [33, 31, 32, 53, 9]. However, uniformly frame sampling may overlook important frames, especially when keyframes of interests are not evenly distributed throughout the video. Compared to existing studies, we propose a vision LLM framework empowered by an adaptive, text-conditioned frame selection mechanism that dynamically selects the most relevant frames for the given vision task.
2.2 Token Merging Techniques
Token merging [7] is a technique that combines similar tokens to reduce computational costs with minimum impact on performance in vision tasks. Concurrently, SD-TOME [6] significantly reduces the computation load of text-to-image generation models through token merging. VID-TOME [28] demonstrates that token merging is also effective for processing visual tokens in videos. In LLMs [21], token merging significantly improves the inference speed by reducing the number of textual tokens. In VLLMs, TOME-LLaVA [46] merges visual tokens to lower computational cost. More recently, Chat-Uni-VI [22] integrates token merging into video LLM pipeline to reduce the number of visual tokens for long videos. LongVLM [54] divides long videos into segments and applies token merging within each segment’s visual tokens. Despite these studies integrating token merging into VLLMs, they often lack flexibility in controlling the number of tokens. Our study for the first time proposes an iterative token merging strategy to achieve finer control over the number of tokens.
Figure 2: Overview of B-VLLM. The frames of the input video are encoded into preliminary visual tokens $V$ including [CLS] tokens $V_{[CLS]}$ by the Vision Encoder. Subsequently, temporal frame selection and token merging use $V_{[CLS]}$ and the text prompt to identity the visual tokens $V^{*}$ corresponding to the frames most relevant to the text prompt. Finally, via the spatial sampling and an optional spatial merging strategy, visual tokens $V^{**}$ sampled from $V^{*}$ are fed into the LLM.
3 Balanced-VLLM
In this section, without loss of generality, we assume that the input to Balanced-VLLM (B-VLLM) is a video-text pair. The video consists of $L$ frames, denoted as $I=\{I_{1},...,I_{L}\}$, where $I_{l}$ represents the $l$-th frame. The text context input, consisting of $N$ words is denoted as $T=\{T_{1},...,T_{N}\}$, where $T_{n}$ is the $n$-th word. The goal of B-VLLM is to control the number of visual tokens when processing long videos, thereby minimizing the risk of exceeding the context window limit of the backbone LLM. The overall architecture of the proposed method is described in Sec. 3.1, followed by the text-conditioned frame selection module in Sec. 3.2, temporal frame token merging in Sec. 3.3, spatial visual token sampling and an optional spatial visual token merging in Sec. 3.4.
3.1 Architecture Overview
The model architecture of B-VLLM is depicted in Figure 2. B-VLLM processes the visual tokens of the input video to keep their number below a pre-set threshold $\theta$. Initially, each frame $I_{l}$ is encoded as a sequence of preliminary visual tokens $V_{l}=\{V_{l,CLS};V_{l,1},...,V_{l,M}\}$ by a vision transformer, where $V_{l,CLS}$ is a [CLS] token that formulates the coarse semantic of the $l$-th frame, $V_{l,j}$ is the $j$-th visual token of $l$-th frame and $M$ is the total number of preliminary frame-level visual tokens except the [CLS] token. In addition, we denote $V=\{V_{1},...,V_{L}\}$ to represent the input video with the visual tokens of each frame. Unlike previous works [33, 30, 9] that discard [CLS] tokens, we explore using [CLS] tokens in video understanding with LLMs.
First, our text-conditioned frame selection module utilizes the coarse semantics captured by $L$ [CLS] tokens in the video and the given text context $T$ to select the $L^{*}$ most informative frames. The selected frames with their visual tokens can be denoted as: $V^{*}=\{V^{*}_{1},..,V^{*}_{L^{*}}\}\subseteq V$. To further lower the total number of visual tokens, we utilize a frame-wise token selection module to sample the $Z$ most informative visual tokens from each $V^{*}_{l^{*}}\in V^{*}$, conditioning on $T$. An optional spatial token merging module is proposed to combine the selected visual tokens that bear high similarity for a further control in the number of tokens. The final selected visual tokens are projected into the LLM feature space and combined with the text context $T$ to form as the input of LLM for downstream tasks, such as video captioning and visual question answering.
3.2 Text-Conditioned Adaptive Frame Selection
Using only the frames most relevant to the given text context is essential for reducing the number of visual tokens. However, identifying these frames can be computationally expensive. To address this, we use the [CLS] token $V_{l,[CLS]}$ for each frame instead of including all other preliminary frame-level visual tokens $V_{l,1},...V_{l,N}$. This design choice balances computational efficiency with the effectiveness of locating relevant frames. We adapt a Q-Former [26] as our frame selection network due to its lightweight design and multimodal reasoning capability.
In detail, the Q-Former creates $L^{*}$ queries $Q=\{q_{1},...,q_{L^{*}}\}$ via jointly encoding the [CLS] tokens $V_{[CLS]}=\{V_{1,[CLS]},...V_{L,[CLS]}\}$ and the text context $T$. These queries are for locating the $L^{*}$ most informative frames from $L^{*}$ perspectives and their corresponding visual tokens. Mathematically, we have:
$V^{*}=S_{\tau}\cdot V=\Phi(Q\cdot V_{[CLS]}^{\top},\ \tau)\cdot V,$ (1)
where $\Phi$ is Gumbel-Softmax operation and $\tau$ is a temperature parameter for the Gumbel-Max trick. Note that $S_{\tau}=\{s_{\tau,l^{*},l}\}$ is an $L^{*}\times L$ matrix, in which $s_{\tau,l^{*},l}$ reflects the weight of $l$-th frame for the $l^{*}$-th selection perspective. As a result, $V^{*}$ is a linear weighted combination of $V$ regarding the weight $S_{\tau}$ for $L^{*}$ perspectives.
Notably, as $\tau$ approaches 0, the Gumbel-Softamx distribution converges to a categorical distribution, meaning that $S_{\tau=0}$ is a matrix where each row is a one-hot vector. Leveraging the property of the Gumbel-Softmax, we construct $\bar{V}$ to enable discretely sampling from $V$. Hence, conceptually, we have $V^{*}_{l^{*}}\in V^{*}$. In particular, if $s_{\tau=0,l^{*},l}=1$, this indicates that ${V}_{l^{*}}^{*}=V_{l}$, meaning $V_{l}$ is selected for the $l^{*}$-thth perspective. In this work, we typically set $\tau$ to 0.1 to ensure the a close approximation of one-hot vectors. Lastly, we restore the temporal order of $V^{*}$ based on $S_{\tau}$ as Gumbel-Softmax may break the order of selected frames. We argue that Softmax is unsuitable in this context as $S_{\tau}$ is calculated based on $V_{[CLS]}$ where each $V_{[CLS],l}$ reflects coarse semantic and may overlook fine-grained spatial details. As a result, Softmax may smooth out rich spatial visual cues especially when aggregating multiple $V_{l}$s.
3.2.1 Discussion: Frame Selection with [CLS] Token.
The role of [CLS] tokens, which carry high-level semantic information of corresponding images [52, 63], has been widely studied, and our ablation studies in Section 5.3 validate their effectiveness in frame selection.
3.3 Temporal Frame Token Merging
The text-conditioned adaptive frames selection retrieves the $L^{*}$ frames with their preliminary visual tokens that are most relevant to the context $T$. However, duplications in the selected frames may occur, for example, when the number of frames $L$ is less than $L^{*}$, leading to duplications. To address this, we propose a temporal visual token merging mechanism to combine similar or duplicated frames selected.
Starting with duplication detection, we use $S_{\tau}$ to identify duplicates in the selected frames. As each perspective (i.e., row-vector) in $S_{\tau}$ approximates a one-hot vector distribution, thereby we can measure the row-wise similarity to locate the potential duplications. With a cosine similarity measure, two row vectors are marked as duplicates if their similarity exceeds a pre-set threshold $\gamma$. We conduct this procedure in an iterative manner. For $V^{*}_{\alpha}\in V^{*}$, we identify and update the frames iteratively as a set:
$D_{\alpha}=\{\beta|\cos(s_{\alpha},s_{\beta})\geq\gamma,\forall s_{\beta}\in S\}\cup\{\alpha\},$ (2)
where $s_{\alpha}$ and $s_{\beta}$ are the $\alpha$-th and the $\beta$-th row vectors in $S$, respectively. Next, we update $V^{*}$ as follows:
$V^{*}\leftarrow V^{*}\setminus\{V^{*}_{\beta}|\beta\in D_{\alpha}\}.$ (3)
The visual tokens of the frames which have indices in $D_{\alpha}$ are merged as follows:
$V^{*}_{\alpha}=\frac{1}{|D_{\alpha}|}\sum_{\beta\in D_{\alpha}}{V^{*}_{\beta}}.$ (4)
Lastly, we append the merged visual tokens back to $V^{*}$ as:
$V^{*}\leftarrow V^{*}\cup\{V^{*}_{\alpha}\}.$ (5)
3.4 Spatial Visual Token Sampling
Although the maximum cardinality of $V^{*}$ is limited, each $V^{*}_{i}$ contains $M$ visual tokens except the [CLS] token, resulting in a total of up to $L^{*}M$ visual tokens. Existing studies [33, 59] typically set $M$ to 256 or 576. When 32 frames are selected, this result in 8,192 or 18,432 visual tokens, respectively, increasing the risk of exceeding the LLM context window limit and imposing a heavy computation load. To reduce the number of visual tokens per frame, we use a spatial Q-Former to sample $R$ visual tokens that are most relevant to the given context $T$, with $R\ll M$. In detail, for each $V^{*}_{l^{*}}$, the spatial Q-Former generates $R$ learnable spatial visual tokens: $V^{**}_{l^{*}}=\{V^{**}_{l^{*},1},...,V^{**}_{l^{*},R}\}$ with the inputs $V^{*}_{l^{*}}=\{V^{*}_{l^{*},1},...,V^{*}_{l^{*},M}\}$ and the text context $T$. Finally, each video contains a total of $MR$ visual tokens.
Notably, the spatial Q-Former generates a fixed number of tokens, which limits flexibility when finer control over the token count is needed. To address this, we propose an optional progressive spatial visual token merging strategy for the scenarios where a maximum visual token count $\theta$ applies. Simply put, if the size of $V^{**}_{l^{*}}$ exceeds the maximum token limit, we halve $V^{**}_{l*}$ by merging the most similar tokens through Bipartite Merging [7]. This process is repeated until the number of tokens is satisfied.
3.5 Integration with Backbone LLMs
As the backbone LLM is a unimodal model that only responds to textual input, we bridge the modality gap between textual and the visual tokens via a learnable MLP model. Next, the projected visual tokens are concatenated with the textual token sequence before being fed to the backbone LLM. Notably, our proposed method is flexible in either integrating with existing VLLM or in being a VLLM on its own. In particular, it is compatible with most of the existing VLLMs architectures, such as LLaMA-VID [30] and VideoLLaMA2 [9]. Here, we denote our method as B-VLLM when there is no integration with other VLLM. In such case, we utilize Qwen2 [56] as backbone LLM for B-VLLM.
 Method #F MVB Perc VNB EgoS VMMEs VMMEm VMMEl MSVDqa MSRVTTqa AcNetqa
Video-LLaVA 8 41.0 44.3 - 38.4 45.3 38.0 36.2 70.7/3.9 59.2/3.5 45.3/3.3
Chat-UniVi-V1.5 8 - - - - 45.7 40.3 35.8 69.3/3.7 55.0/3.1 46.1/3.3
VideoChat2 16 51.1 48.3 12.4 33.2 48.3 37.0 33.2 70.0/3.9 54.1/3.3 49.1/3.3
Sharegpt4Video 16 51.2 - - - 48.3 36.3 35.0 45.6/- 43.0/- 50.8/-
LLaVA-NeXT-V 32 46.5 48.8 - 43.9 36.0 32.9 28.0 67.8/3.5 -/- 53.8/3.2
LLaMA-VID 1fps 39.0 41.2 5.0 31.2 34.2 34.7 27.1 69.7/3.7 57.7/3.2 42.4/3.2
w. Ours 1fps 43.5 47.0 11.6 38.2 44.7 38.8 35.2 71.5/4.0 57.9/3.5 43.9/3.4
[…]
Video-LLaMA2 8 45.5 45.6 10.8 42.2 48.9 42.7 37.7 70.2/3.9 54.8/3.3 47.6/3.4
w. Ours 1fps 46.5 48.0 12.6 44.3 47.2 44.4 41.5 72.1/4.0 58.1/3.5 48.5/3.4
[…]
B-VLLM 1fps 50.8 52.1 22.6 51.9 60.8 51.8 47.9 72.3/4.0 59.2/3.5 48.6/3.7
Table 1: Quantitative comparisons on video benchmarks in terms of accuracy $\uparrow$.
4 Implementation
Baselines. We evaluate against two baselines: LLaMA-VID [30], which uses high spatial compression for video understanding, and VideoLLaMA2 [9], which uniformly samples 8 frames per video. Datasets & Training. B-VLLM is trained solely on LLaMA-VID-Dataset [30] and Valley [38] for fair comparison. See supplementary material for details.
 Method # Token. MMB POPE VizWiz MME VQAv2 SciQA GQA TQA SEED
InstructBLIP 32 36 - 34.5 - - 60.5 49.2 50.1 58.8
Qwen-VL 256 38.7 - 35.2 - 78.8 67.1 59.3 63.8 56.3
Qwen-VL-Chat 256 60.6 - 38.9 1488 78.2 68.2 57.3 61.5 58.2
LLaVA-1.5 576 64.7 85.6 50.0 1510 78.5 66.8 62.0 58.2 66.1
LLaMA-VID 2 49.8 75.3 37 1187 62.7 65.3 48.8 42.1 50.8
w. Ours 32 59.3 83.8 36.9 1285 73.0 66.9 53.0 46.6 61.0
[…]
VideoLLaMA2 72 64.6 84.3 46.7 1455 74.5 72.2 56.9 54.6 63.9
w.Ours 32 67.3 85.3 47.3 1501 74.5 72.0 59.7 52.8 66.7
[…]
B-VLLM 32 71.9 84.5 45.1 1455 76.6 74.5 58.4 54.8 67.1
Table 2: Quantitative comparisons on image benchmarks in terms of accuracy $\uparrow$.
5 Experimental Result & Discussion
5.1 Video Benchmark Quantitative Evaluation
We evaluate the proposed B-VLLM on various video benchmarks in a zero-shot manner. The video duration in the these benchmarks spans from 10 second to 1 hour, where the details are summarized in the supplementary materials. For open-end answering, we have MSVD-QA [55], MSRVTT-QA [55] and ActivityNet-QA [12], which aim to evaluate the capacity of video captioning. For multi-choice question answering (MCQA), we have MVBench [27] and VideoMME [15], which are multi-purpose video understanding benchmarks; EgoSchema [40] focuses on evaluating the temporal understanding capability; Perception [43] aims to examine reasoning and perception skills of VLLMs; and VNBench [60] evaluates a VLLM using the needle in a haystack manner: asking questions about one or more irrelevant images that are randomly inserted into a video.
Quantitative results for open-end question answering and MCQA benchmarks are reported in Table 1. Overall, B-VLLM achieves SOTA performance on MSVD-QA, Perception, VNBench, Egoschema, and VideoMME. Specifically, when integrating with existing VLLMs, it significantly boosts performance to the two baselines across most video benchmarks. Notably, our method achieves substantial improvements in long-video scenarios. For example, VideoMME-Long evaluates VLLMs performance in extremely long-video settings, with an average video duration of 2466.7 seconds. Our proposed method introduced an 10.5% performance gain for LLaMA-VID [30] and 3.8% for VideoLLaMA2 [9]. Context window length is one of the key factors limiting existing VLMs to handle long videos. For example, LLaMA-VID encodes each frame with 2 tokens, it struggles with extremely long video due to its context window limit and potential key information loss. Our adaptive frame selection module addresses this limitation by selecting a fixed number of frames most relevant to the given context, ensuring the number of visual tokens remains within the context window of LLaMA-VID. As a result, LLaMA-VID consistently performs well in medium and long video settings. We also observe a noticeable performance drop when VideoLLaMA2 handles medium and long videos, which we attribute to its uniform frame sampling strategy that may overlook the frames relevant to the context. The integration of our method mitigates this issue by filtering out irrelevant frames, leading to 1.7% and 3.8% performance gains for VideoLLaMA2.
Additionally, the proposed method remains effective in the short-video settings. For example, LLaMA-VID achieves a 4.5% and 5.8% performance gain in MVBench and Perception through the integration of our method. Similarly, in open-end benchmarks, performance improvements of 1.7%, 0.2% and 1.5% are observed in MSVD-QA, MSRVTT-QA and ActivityNet-QA, respectively. This substantial improvement can be attributed to the enhanced spatial information from the increased number of spatial visual tokens compared to LLaMA-VID. Similar performance improvements are also observed in Video-LLaMA2, with gains of 1.0%, 2.4%, 1.8%, 2.1%, 1.9%, 3.3% and 0.9% on MVBench, Perception, VNBench, Egoschema, MSVD-QA, MSRVTT-QA and ActivtiyNet-QA, respectively.
An interesting observation is that VLLMs employing uniform sampling achieve better performance in short-video scenarios, particularly when the number of sampled frames closely matches the video length in seconds. For example, VideoChat2 [27] ranks highest on MVBench (16 frames), and LLaVA-Next-Video [59] ranks second on Perception (23 frames). Although these methods could theoretically improve performance in longer videos by sampling more frames, they quickly exceed the context-window limitations, increasing computational costs and reducing performance. In contrast, our method maintains consistent performance across short and long videos due to effective frame selection and spatio-temporal merging techniques. Refer to the supplementary material for more experimental result.
5.2 Image Benchmark Quantitative Evaluation
Following [33], we select MMBench [34], POPE [29], VizWiz [17], MME [14], VQAv2 [16], ScienceQA-Image [36], GQA [20], TextQA [49] and Seed-Image [25] to evaluate the generalizability of B-VLLM in the image setting. Evaluation results are reported in Table 2.
Our proposed method yields comparable, if not superior, performance, despite not being specifically designed for the image-only setting and using significantly fewer visual tokens per image than existing image-based VLLMs. In particular, B-VLLM outperforms classic image-only VLLM, LLaVA-1.5 [33] on MMBench [34], ScienceQA-Image [36], and SEED-Image [25] by 7.1%, 7.7%, and 1.0%, respectively. Notably, comparable results are observed on POPE [29] and MME [13], despite LLaVA-1.5 [33] consuming 18 times more visual tokens than ours. These benchmarks primarily focus on reasoning and general image understanding, where fine-grained visual details are less critical and typically require fewer visual tokens. However, as noted by LLaMA-VID [30], more visual tokens generally lead to better performance on certain image-only benchmarks. The limitation in the number of visual tokens compromises our method’s performance on VizWiz [17], VQAv2 [16] and TextVQA [49], where these benchmarks examine spatial reasoning and OCR. For example, TextVQA [49] focuses on detecting text within image, while GQA [20] emphasizes complex scene understanding.
In terms of performance gain, integrating our proposed method results in significant improvements for LLaMA-VID [30] and VideoLLaMA2 [9] on image benchmarks. For LLaMA-VID, we observe gains of 9.5% and 10.2% on MMB and VQAV2; and for VideoLLaMA2, non-trivial gains can be observed when incorporating with our method. We believe these gains stem from the effectiveness of the visual token sampling module, as it dynamically selects the most informative tokens, reducing the overall token count while maintaining performance.
 F.S. T.M. $S.$ MVB VMME MMB POPE
39.0 33.6 49.8 75.3
39.4 36.5 58.0 81.1
42.1 38.1 54.6 67.3
43.5 39.6 59.3 83.8
Table 3: Ablation study on adaptive frame selection (F.S), temporal frame token merging (T.M.), and spatial visual token sampling (S.S). Experiments are done on LLaMA-VID.
(a) Impacts on the number of frames selected LL* and spatial vis. tokens $R$
(b) Impacts on the temperature $\tau$ and frame similarity threshold $\gamma$
(c) Impacts on maximum number of token $\theta$
Figure 3: Impacts on the hyperparameters.
Figure 4: Qualitative examples for two short-medium videos and one long video.
5.3 Ablation Study
Effectiveness of the proposed modules in terms of their performance is reported in Table 3. In this experiment, LLaMA-VID [30] is selected as the base model. MVBench [27] and VideoMME [15] are used to evaluate video understanding performance, while MMBench [34] and POPE [29] are used to assess image understanding performance. Starting with the baseline, without the integration of our proposed modules, LLaMA-VID [30] displays moderate performance on the short-video benchmark MVBench [27]. However, compromised performance can be observed on the medium-to-long video benchmark Video-MME [15] due to the limitation of the context window length. In addition, LLaMA-VID [30] typically performs worst on image benchmarks, due to its poor spatial perception caused by a deficit of spatial visual tokens. With the introduction of the text-conditioned adaptive frame selection (F.S.) module, the baseline model no longer suffers from its own token selection mechanism, resulting in 2.9% performance boost on the medium-to-long video benchmark VideoMME [15]. With the temporal frame token merging (T.M.) mechanism, the model mitigates temporal redundancy in duplicated frame selection, resulting in noticeable performance gains on both short and medium-to-long video benchmarks. Similarly, the integration of spatial visual token sampling (S.S.) enables dynamic adjustment on the number of spatial visual tokens to balance the temporal and spatial cues, resulting in performance gains 1.4% and 1.5% on MVBench and VideoMME, respectively. More ablation experiments can be found in supplementary material.
Ach./Token. Time MVB EgoS Perp. VMME
Random - 27.7 20.0 33.0 25.0
Q-Former 25 hrs. 46.5 44.3 48.0 44.5
Resampler 23 hrs. 44.3 45.5 46.8 44.0
Mean Pooling 19 hrs. 41.2 40.0 41.1 37.6
CLS Token 10 hrs. 34.5 34.1 37.8 36.8
Table 4: Q-Former v.s Resampler in Frame Selection and Mean Pooling v.s [CLS] tokens in VLLMs.
Impacts on the number of frame selected LL* , spatial visual tokens $R$ and maximum visual token count $\theta$ are illustrated in Figure 3(a) and  3(c). B-VLLM is selected as base model in the upcoming experiments.First, we examine the impact of two variables: the number of selected frames LL* and the number spatial visual tokens $R$ per frame. More specifically, we investigate how different combinations of these variables affect the performance of a trained model. Interestingly, as shown in Figure 3(a), we observed that, compared to the maximum setting with 32 selected frames and 32 spatial visual tokens per frame, it exhibits only a marginal performance drop when the number of selected frames is reduced while maintaining a sufficient number of visual tokens. For example, when the number of spatial visual token $R$ is set to 32, reducing the number of selected frames LL* from 32 to 16 results in a performance drop of 0.9% and 1.3% on MVBench and VideoMME-Medium, respectively. In a more extreme setting, reducing the number of frame selected LL* from 16 to 8 results in a performance drop of 0.5% and 0.4% performance drop on MVBench and VideoMME-Medium, respectively. We argue that this is because 8 to 32 key frames are sufficient for most video understanding tasks. When using 32 frames, dropping the number of visual tokens $R$ from 32 to 16 results in performance drops of 0.6% on MVBench and 0.4% on VideoMME-Medium. This indicates that current video understanding tasks tend to rely more on temporal dynamics rather than on detailed spatial information. Secondly, we examine how the maximum token count $\theta$ affects the trained model while keeping both the number of selected frames LL* and the number of spatial visual tokens $R$ fixed at 32. As shown in Figure 3(c), increasing the token count generally improves performance, but the gains for B-VLLM plateau at 512 tokens. This indicates B-VLLM can deliver strong performance without a high computational budget.
Impacts on temperature $\tau$, similarity threshold $\gamma$ are presented in Figure 3(b). In these experiments, the number of selected frames LL* and spatial visual tokens $R$ are both set to 32, and the maximum token count $\theta$ is fixed at 768. Our results indicate that the optimal performance is achieved with a lower $\tau$ value in combination with a higher $\gamma$ value. Where $\tau$ governs the discreteness of Gumbel-Softmax during frame selection and higher $\gamma$ values prevents dissimilar frames from being merged. For example, when $\tau$ is set to 0.1, reducing $\gamma$ from 0.75 to 0.5 results in 1.2% and 1.7% performance drop on MVBench and VideoMME-Medium, respectively. Similarly, when $\gamma$ is fixed to 1.0, lowering $\tau$ from 0.1 to 0.5 results in 0.7% and 1.2% performance drops. We find that a high $\gamma$ prevents merging dissimilar frames, while a low $\tau$ mitigates the oversmoothing of frame features caused by the Gumbel-Softmax.
Q-Former v.s. Resampler in Frame Selection. Besides Q-Former, we also verified with other commonly used frame selection module - Resampler, where the results are reported in Table 4. Resampler is originally designed for visual token abstraction, for fair comparison, we provide it with both visual and text tokens, setting the same number of query tokens as in Q-Former. Notably, Q-Former outperforms Resampler on most selected benchmarks. We argue that the Resampler’s reliance on self-attention mechanism limits its performance in processing cross-modality tokens.
Efficacy of [CLS] token. To validate the efficacy of [CLS] tokens in frame selection, we compare its performance against using visual tokens with mean pooling as frame features. For simplicity, we conduct preliminary experiment on VideoLLaMA2 [9] where the modality projector is replaced with a single MLP layer. It is worth mentioning that, none of the proposed module is present in this preliminary experiment. All experiments are trained using the same settings, with 8 frames selected from each video. As shown in Table 4, [CLS] tokens enables the model to consistently outperform random guessing while being more training-efficient than mean pooling. This efficiency is preferable and the effectiveness is sufficient for frame selection, as the goal is to filter out most irrelevant frames in a coarse stage. Refer to the supplementary material for more details.
5.4 Discussion
Visualization. As shown in Figure 4, we visualize frame selection and spatial tokens for two short-medium videos and one long video. In the first case, the frame selection module locates the frames with the banana taped to the wall. Then based on the question, the visual tokens near the banana and tape are sampled and used by LLM, Lastly, the LLM backbone answers the question correctly based on the selected visual tokens. In the second video, the selected frames are from both ends, demonstrating question-based selection rather than random or uniform sampling. In contrast, for long videos with overarching questions, the frames selected are evenly distributed. Limitation. We recognize that using [CLS] tokens for frame selection has its limitations since pooled visual tokens are more informative, as an alternative strategy for exploring more accurate frame selection. See the supplementary material for further discussion.
6 Conclusion
We present a novel vision LLM framework - B-VLLM, designed to address the issue of visual token overload in video understanding with LLMs. Our approach controls the number of visual tokens by adaptively selecting the most relevant frames and applying dynamic spatial token sampling and merging. Comprehensive experiments confirms the effectiveness of B-VLLM. We also highlight the limitations of our method to inform future directions.
Acknowledgements
This work was supported by the Australian Research Council (ARC) Linkage Project #LP230100294, ARC Discovery Project #DP210102674 and Edith Cowan University Science Early Career and New Staff Grant Scheme.
Supplementary Material
Table 5: Statistics of training datasets.
 Stage Modality Sample Source
LLaMA-VID [30]
PT Image 558K  [47, 33]
Video 232K [5, 30]
  FT Image 625K  [16, 20, 23, 24, 41, 42, 45, 48]
Video 98K  [12]
Text 40K  [1]
VideoLLaMA2 [9] - Valley [38]
PT Image 558K  [47, 33]
Video 702K [38]
  FT Image 665K  [33]
Video 100k  [39, 38]
Table 6: Statistics of benchmark datasets. Type indicates the QA question type: Open represents open-ended answering and MCQA denotes multi-choice question answering.
 Benchmark Avg. Len (s) # QA Pair Type
MSVD-QA 9.8 10K Open
MSRVTT-QA 15.2 50.5K Open
MVBench 16 4K MCQA
Perception-Test 23 2.7K MCQA
Next-QA 44 4000 MCQA
VNBench 54 5.6K MCQA
VideoMME-S 80.7 0.9K MCQA
ActivityNet-QA 111 8K Open
EgoSchema 180 5K MCQA
VideoMME-M 515.9 0.9K MCQA
VideoMME-L 2466.7 2.7K MCQA
7 Additional Implementation Details
7.1 Training Details
We adopt a two-stage training strategy [33, 30, 9], dividing training into pretraining for modality alignment and fine-tuning for instruction tuning. During pretraining, the learning rate is set to 1e-3 with a linear learning rate schedule, and the total batch size is set to 256. Only the frame selection module and the MLP projection layer, which maps the visual token features to the LLM features, are updated. In the fine-tuning stage, the entire framework is unfrozen except for the preliminary visual encoder. Additionally, LoRA [18] is enabled with the rank set to 128 and $\alpha$ set to 256. The total batch size was set to 128, and the learning rate is 1e-4 with cosine scheduling. Supervision is provided by minimizing cross-entropy for masked text tokens, and the optimizer for both stages is AdamW [35], with DeepSpeed ZeRO2 [3] enabled for memory efficiency. Training is conducted with 16 Nvidia L40S GPUs. Following previous works [27, 39, 9], videos are downsampled to 1 fps in both training and evaluation.
7.2 Dataset for Training
The following provides details on the training datasets utilized in LLaMA-VID-Dataset [30] and VideoLLaMA2 [9]. The dataset used by LLaMA-VID [30] consists of 790K samples for pre-training stage and 763K samples for fine-tuning stage. Compared to the dataset used by LLaMA-VID, Valley [38] contains significantly more video samples for pre-training. There are 1.25M samples for pre-training stage and 765K sample for fine-tuning stage. The difference in the volume of video samples may explain the performance differences observed between LLaMA-VID and VideoLLaMA2.
7.3 Video Benchmark Summary
The details of the video benchmarks are summarised in Table 6. We selected benchmarks with diverse durations, avoiding a focus on a fixed average video length. For example, the MSVD-QA [55] benchmark consists of short videos with an average duration of 9.8 seconds. In contrast, the VideoMME-L [15] is a representative of long video benchmark, with an average duration of 2466.7 seconds and a maximum video length capped at 1 hour. We argue that our selection of video benchmarks reflects the diversity of video durations in real-world scenarios, demonstrating the applicability and robustness of our model in practical environments.
7.4 B-VLLM Experimental Setup
Our B-VLLM is implemented based on VideoLLaMA2 [9], where Qwen2 [56] is selected as the backbone LLM. Each frame is encoded as 768 preliminary visual tokens using EVA-CLIP [sun2023eva], Subsequently, 32 frames are selected using the frame selection module and temporal frame token merging module. The visual tokens corresponding to the selected frames are encoded into 32 spatial visual tokens by our spatial visual token sampling module.
8 Additional Experiment & Discussion
8.1 Additional Discussion on Different Frame Selection Features.
As reported in Table 7, we additionally report the performance of selecting frames by using three other feature extraction methods: (a) max pooling (b) mean pooling and (c) Qformer. To investigate the impact of these variants in B-VLLM’s fine-grained spatial perception, we also report the performance of OCR and Counting in VMME, which are more relevant to spatial capability. Though [CLS] token may not be the optimal choice in terms of performance, as discussed in Section 3.2, the rationale of using [CLS] token lies in balancing computational cost and effectiveness.
Table 7: Additional ablation study on frame selection features.
Model Train. MVB EgoS VMME OCR Counting
Mean 13.1h 51.3 51.9 54.4 48.2 36.6
Max 13.1h 49.5 51.6 50.7 45.3 38.1
Qformer 14.0h 52.4 52.9 51.6 41.7 35.8
CLS 10.9h 50.8 51.9 52.9 46.0 34.0
8.2 The Impact of LL*.
As shown in Fig 5, we report additional experiments to investigate the impact of LL* on Short ($\leq$ 3min), Medium ($\leq$ 15min) and Long video ($\leq$ 60 min) settings. As shown below, B-VLLM’s performance generally saturated at L=28L*=28, though it was trained with L=32L*=32. This indicates that there could be an optimal setting for LL*. We argue that this is due to most tasks typically rely on a small portion of frame regardless of the duration of the video. However, we acknowledge that increasing LL* is potentially beneficial for more extreme setting and this review inspired us to develop a LL*-free design as one of our future work.
Figure 5: Additional experiment on the impact of LL*.
8.3 Model’s Parameter Size Comparison
In Table 8, we report our B-VLLM’s parameter size below along with TFLOPs and GAMCs, which are similar to most existing methods.
Table 8: Model size compared with existing methods
Model TFLOPs GMACs Param.
InternVL2 1.83 910 7+1.1B
InternVideo2 1.83 915 7+1.5B
QwenVL 1.92 915 7.7+1.9B
Qwen2VL 1.82 905 7.6+0.6B
VideoLLaMA2 1.82 910 7.6+0.8B
B-VLLM (Ours) 1.81 908 7.6+1.2B
Table 9: Additional Comparison with Recent SOTA methods
Model #Train.Data MVB EgoS Perp. VMME
InternVideo2 39M 60.3 55.8 53.0 41.9
InternVL2 20M+ 66.4 - - 54.0
LLaVA-OV 9.3M 56.7 60.1 57.1 58.2
Qwen2VL - 67.0 66.7 62.3 66.3
B-VLLM (Ours) 2M 50.8 51.9 52.1 52.9
8.4 Additional Comparisons with Recent SoTA.
We provide additional comparisons with several advanced VLLMs in Table 9. Note that these advanced VLLMs are generally trained with at least 5 times more data than our B-VLLM, and these data are often private. This indicates the great potential of our B-VLLM when it is trained with large scale data.
8.5 Inference Speed & Latency
In Table 10, we compare the inference speed between Qwen2VL and B-VLLM on three types of videos: 4-minute video, 8-minute video, and 12-minute video. As reported below, B-VLLM demonstrates clear advantage on inference speed. Specifically, it takes 5.6s for B-VLLM to generate a response for an 8-minute video while taking 24.6s for Qwen2-VL to respond.
Table 10: Inference Speed and Latency
Model 4 Min 8 Min 12 Min
Qwen2VL 6.94s 24.6s 26.2s
B-VLLM (Ours) 3.16s 5.6s 7.9s
8.6 Additional Discussion on Limitations
Limitation. On top of the discussed limitation in the main manuscript, we acknowledge that our proposed method has limitations in handling multi-round conversations based on the same video. In the video-based multi-round conversation, B-VLLM repeats the frame selection and spatial visual token procedure as the context shifts during conversation, resulting in additional computation cost. Moreover, as discussed in manuscript, the spatial visual tokens in B-VLLM are underutilized, and the full potential of our method for image data has yet to be fully explored. Additionally, the Gumbel-Softmax sampling technique may disrupt the temporal order of selected frames, potentially compromising our method’s performance in temporal perception. Addressing these limitations could further enhance the capabilities of VLLMs. We recognize that using [CLS] tokens for frame selection has its limitations since pooled visual tokens are more informative, as an alternative strategy for exploring more accurate frame selection. More limitations are discussed in supplementary material. Additionally, the fixed value $L^{*}$ can limit B-VLLM’s performance on extreme long videos, especially when the number of relevant frames surpasses $L^{*}$.
9 More B-VLLM Qualitative Examples
This section provides additional conversation examples with B-VLLM based on videos. Note that the presented videos are accessible only through the specified source. The examples are depicted in Figures 6 to 10. As shown in Figure 6, when a game trailer is fed into B-VLLM, it accurately identifies the game’s title displayed at the beginning of the video and effectively summarizes the gameplay content, demonstrating B-VLLM’s comprehensive understanding of both temporal and spatial information. In the example shown in Figure 9, during the conversation, B-VLLM accurately identifies the man’s actions. Even though the basketball appears only briefly at the beginning of the video, B-VLLM successfully captures and localizes the nuanced spatial details. As depicted in Figure 10, beyond accurately describing the content of the given video, B-VLLM demonstrates the ability to estimate the age of the man featured in the video, highlighting the spatial perception capabilities of our method.
Figure 6: Visualization on understanding a 2-minutes game trailer containing intensive fighting scenes.
Figure 7: Visualization on understanding a 1-minute AI generated video of a woman in Japan.
Figure 8: Visualization on perceiving a 4-minute wonderful music video containing scenes of singing and dancing.
Figure 9: Visualization on perceiving a 1-minute dancing video a man showcasing his outstanding dancing and basketball skills.
Figure 10: Visualization of understanding a 3-minute video featuring a Kung Fu master demonstrating his martial arts expertise.